OpenStudy (anonymous):

A coffee mixture has beans that sell for $1.35 per pound and beans that sell for $1.05 per pound. If 120 pounds of beans create a mixture worth $1.15 per pound, how many pounds of each bean is used?

OpenStudy (anonymous):

($/lb)*lb = $ Let the required weight in lb of the $1.35 per pound beans be x. Solve the following equation for x: x (1.35) + (120 - x)(1.05) = 120(1.15)